OpenStudy (anonymous):

Solve the DE: y'+(2x-1)y=4x-2

OpenStudy (anonymous):

I got the answer \[y=2+\frac{ce^x}{e^{x^2}}\]

OpenStudy (anonymous):

it can be simplified more i guess

OpenStudy (experimentx):

\[ \huge y = e^{-\int (2x-1)dx} \times \int (4x-2)e^{\int (2x-1)dx} dx + \\ \huge Ce^{-\int (2x-1)dx}\]

OpenStudy (anonymous):

that's what i did where'd the - come from though

OpenStudy (anonymous):

for the first integral and last?

OpenStudy (anonymous):

the integrals equal \[e^{x^2-x}\]

OpenStudy (anonymous):

i can't remember the property of base e is it \[e^{a^b}=e^{a*b}\]

OpenStudy (anonymous):

if that is the case then it can be simplified more

OpenStudy (experimentx):

Oh ... there's a whole lecture series discussing how to solve this type of equation here http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/

OpenStudy (anonymous):

did you ever get an answer? that was my personal answer... i needed a check lol

OpenStudy (anonymous):

it really isn't as base as it looks because the x^2-x when used as a u will get u du=(2x-1)dx

OpenStudy (anonymous):

bad* as it looks

OpenStudy (experimentx):

seems you are right http://www.wolframalpha.com/input/?i=y%27+%2B+%282x-1%29y+%3D+4x-2

OpenStudy (anonymous):

thanks

OpenStudy (experimentx):

yw